OpenStudy (anonymous):

can anyone help me solve for y? integral(dy/(100-y)) = integral(xdx)

OpenStudy (anonymous):

\[\int\limits_{?}^{?}(1/100-y)= \int\limits_{?}^{?}(x*dx)\]

OpenStudy (turingtest):

\[\int\frac{dy}{100-y}=\int xdx\]\[u=100-y\to du=-dy\]\[-\int\frac{du}{u}=\int xdx\]\[-\ln u=\frac12x^2\]\[\ln(100-y)=-\frac12x^2\]\[100-y=e^{-x^2/2}\]\[y=100-e^{-x^2/2}\]simplify if you wish...

OpenStudy (anonymous):

ok i see thank you

OpenStudy (anonymous):

I believe there should have been an absolute value sign around the u...which would make it + or minus e^...

OpenStudy (anonymous):

But the rest of it looks good.

OpenStudy (anonymous):

don't for get the c at the end.

OpenStudy (anonymous):